OpenStudy (anonymous):

Mrs Brown invested $15,000 in two account, one yielding 7% interest and the other yielding 12%. If she received $1,600 in interest how much did she invest in each account?

OpenStudy (anonymous):

in the account that has 12%, he invested 11000 and in the other 7% account he invested 4000

OpenStudy (anonymous):

Let x be the amount invested in the first account. Solve the following for x: x (.07) + (15000 - x) (.12) =1600 x = 4000, 15000 - x = 11000 Check the result: 4000 * (.07) + (11000) * (.12) = 1600.